OpenStudy (anonymous):

A satellite moves in a circular orbit around the Earth at a speed of 6.3 km/s. Determine the satellite’s altitude above the surface of the Earth. Assume the Earth is a homogeneous sphere of radius 6370 km and mass 5.98 × 1024 kg. The value of the universal gravitational constant is 6.67259 × 10−11 N · m2/kg2. Answer in units of km

OpenStudy (anonymous):

Well it is a physics problem but I tried using the gravitational constant one I cant figure it out

OpenStudy (isaiah.feynman):

I know its a physics problem, but what topic?

OpenStudy (anonymous):

I'm not sure honestly :/

OpenStudy (isaiah.feynman):

Strange, where did you pick this from then?

OpenStudy (anonymous):

It's from a review for my physics test tomorrow

OpenStudy (isaiah.feynman):

Sure that's all the information?

OpenStudy (anonymous):

That's all it gave me I've been trying to figure out how to get the mass of the satellite but I haven't figure it out since they once gave the speed

OpenStudy (isaiah.feynman):

Give me the link to the place you got this question from.

OpenStudy (anonymous):

Hmmm idt I can because it's through the university and you need a password to get in

OpenStudy (john_es):

You need the formula of orbital velocity, \[v=\sqrt{\frac{GM}{R}}\] with M the mass of the planet and R the radius of the orbit.

OpenStudy (john_es):

The radius of the orbit, also, is, \[R=R_T+h\] where R_T is the radius of the planet and h is the altitude above the surface of the planet.

OpenStudy (john_es):

Can you continue with this information?

OpenStudy (john_es):

You should solve for h, and obtain something like \[h=\frac{GM}{v^2}-R_T\]

OpenStudy (anonymous):

Ok thanks I'll try that :)

OpenStudy (john_es):

;)

OpenStudy (john_es):

You can deduce that you should use the orbital velocity because it is a circular movement, so, the gravitational force is a centripetal force, in modulus, \[G\frac{mM}{R^2}=m\frac{v^2}{R}\]Solving for v, you obtain the orbital velocity.

OpenStudy (anonymous):

but m would be the mass of the satellite?

OpenStudy (john_es):

Exact. The mass of the satellite as you can see can be eliminated, and doesn't matter.

OpenStudy (john_es):

\[G\frac{mM}{R^2}=m\frac{v^2}{R}\Rightarrow G\frac{M}{R^2}=\frac{v^2}{R}\]

OpenStudy (anonymous):

I got 250281 for the velocity so I use this one or the 6.3 km/s for the previous problem?

OpenStudy (john_es):

No, the velocity cannot change, it is v=6.3 Km/s

OpenStudy (john_es):

Use the equation I gave you, \[h=\frac{GM}{v^2}-R_t=\frac{6.67\cdot10^{-11}\cdot5.98\cdot10^{24}}{(6.3\cdot10^3)^2}-6.37\cdot10^6\]

OpenStudy (isaiah.feynman):

@John_ES Escape velocity is \[V_{Escape} = \sqrt{\frac{ 2GM }{ R }}\]

OpenStudy (john_es):

You don't need escape velocity, you need orbital velocity because you the satellite doesn't need to go out the Earth. The satellite is orbiting the Earth so it has orbital velocity.

OpenStudy (isaiah.feynman):

I haven't studied up to that point.

OpenStudy (john_es):

You can check a brief introduction here, http://en.wikipedia.org/wiki/Orbital_speed

OpenStudy (john_es):

The solution I obtain is \[h=10^{13}\ m\] Remember to put all distances in meters.

OpenStudy (anonymous):

Hmmm I got 3.67953*10^6

OpenStudy (john_es):

You're right I don't put the right numbers. ;)

OpenStudy (john_es):

In my calculator XD.